### v1

# Analytical Emulator for the Baryon EoM inside the Fuzzy Dark Matter Soliton from Machine Learning

Abstract
An empirical baryon density profile can be included in the Schrödinger-Poisson (SP) equations to influence the fuzzy dark matter (FDM) soliton formation. However, to probe the effects of baryon on the other dynamical evolutions of the FDM soliton, its equation of motion (EoM) inside the corresponding FDM soliton is needed. In this paper, given an empirical baryon density profile, we first provide the cylinderical symmetric FDM soliton solution about the FDM density and the total potential of FDM and baryon. Then, we build an analytical baryon EoM from the obtained FDM density and total potential by machine learning . Finally, we check that this baryon EoM works as well as an empirical baryon density profile for the FDM soliton formation, with the fractional errors $\lesssim 0.04$. It should also work well for some other simple FDM soliton evolutions.
pacs:
I Introduction
As a promising candidate for dark matter Rubin:1982kyu ; Davis:1985rj ; Clowe:2006eq , the ultralight scalar field with spin-0, extraordinarily light mass ($m\sim 10^{-22}~\rm{eV}/c^{2}$) and de Broglie wavelength comparable to a few kpc, namely fuzzy dark matter (FDM) Hu:2000ke , can form an equilibrium configuration with size smaller than its de Broglie wavelength, the so-called FDM soliton Guzman:2004wj . The subsequent evolution of FDM solitons are usually simulated numerically, including the perturbation, the interference/collision and the tidal disruption/deformation of FDM solitons Guzman:2004wj ; Bernal:2006ci ; Paredes:2015wga ; Edwards:2018ccc ; Munive-Villa:2022nsr ; Tan:2024spp , according to the coupled Schrödinger–Poisson (SP) system of equations. If FDM solitons keep stationary in a fixed background potential due to supermassive black holes Davies:2019wgi , an additional cylindrical Bar:2019bqz or ellipsoidal Tan:2024dne baryon density profile, the shooting method or the successive over-relaxation method can be used to compute the eigenvalues of equilibrium configurations.
However, how to deal with the dynamic evolution of FDM solitons in an evolving background potential? In this case, we should enlarge the SP system with an additional equation of motion (EoM) to describe the evolving background potential. For example, to simulate the FDM soliton collisions in the evolving background potential due to an simultaneously evolving baryon density profile, the baryon EoM inside the FDM soliton should be built according to the interaction between FDM and baryon and the thermodynamical properties of baryon in galaxies. Although some general DM-baryon relations in galaxies have been found, such as a universal constant for DM-baryon interplay Chan:2019ukj , the relations between the total mass of DM and baryon in galaxies Chan:2022fzq , the mass discrepancy-acceleration relation McGaugh:2004aw , the radial acceleration relation McGaugh:2016leg and so on, these relations are too rough to be used to construct the baryon EoM inside the FDM soliton.
In this paper, we propose a novel method to construct the baryon EoM inside the FDM soliton given a mock data of FDM and baryon density profile inside the FDM soliton. Usually, it’s very easy to reconstruct a function of one variable $f(x)$ according to data, such as reconstruction of equation of state of dark energy $w(z)$, Hubble parameter $H(z)$, luminosity distance $D_{L}(z)$ or cosmic reionization history $x_{\rm e}(z)$ by MCMC global fitting Collett:2019hrr , principal components analysis Huang:2017huh , Gaussian process Shafieloo:2012ht ; Seikel:2012uu ; Yang:2024rsy , artificial neural networks(ANNs) Wang:2019vxv or genetic algorithms (GAs) Arjona:2019fwb ; Nesseris:2012tt . As for the function of multiple variables f(x,y,)f(x,y,...), machine learning methods including GAs and ANNs are also the better choices for its reconstruction from data, such as the built of the analytical emulator for the linear matter power spectrum $P(k;\omega_{\rm b},\omega_{\rm m})$ by GAs Orjuela-Quintana:2022nnq ; Orjuela-Quintana:2023uqb ; Orjuela-Quintana:2024hha ; Bartlett:2023cyr and the gravitational wave surrogate models by ANNs Setyawati:2019xzw ; Khan:2020fso ; Thomas:2022rmc . Since the baryon EoM is supposed to be a function of FDM density and total potential $\rho_{\rm b,EoM}(\rho,\Phi)$, we will turn to GAs to build it too.
This paper is organized as follows. In Section II, we solve the SP system with an empirical baryon density profile background in cylindrical coordinates. In Section III, we extract a mock data of baryon density profile and then construct the analytical emulator for baryon EoM by GA. In Section IV, we check our baryon EoM by including it an enlarged SP system. Finally, a brief summary and discussion are provided in Section V.
II FDM soliton solution with an empirical baryon density profile
Given an empirical baryon density profile background, FDM obeys the following SP system,
$\begin{cases}\begin{aligned} &i\hbar\frac{\partial\psi}{\partial t}=\left(-\frac{\hbar^{2}}{2m}\nabla^{2}+m\Phi\right)\psi,\\ &\nabla^{2}\Phi=4\pi G|\psi|^{2}+\rho_{\rm b},\end{aligned}\end{cases}$ (1)
where $m$ is the mass of the FDM particle, $\psi$ is its wavefunction, and $\Phi$ is the total gravitational potential, which is sourced by the FDM density $\rho=|\psi|^{2}$ and the baryon density $\rho_{\rm b}$. For the Milky Way, as Bar:2019bqz ; Launhardt:2002tx , we only consider the main contributions: a nuclear stellar disk (NSD) and a spherical nuclear cluster (NSC)
$\rho_{\rm b}(R,z)$ $=\rho_{\rm NSD}(R,z)+\rho_{\rm NSC}(r),$
$\rho_{\rm NSD}(R,z)$ $=\frac{\bar{\rho}_{\rm NSD}}{1+\left(\frac{R}{250{\rm pc}}\right)^{14}}\left[1-\tanh^{4}\left(\frac{R}{140{\rm pc}}\right)\right]e^{-\frac{|z|}{15{\rm pc}}},$
$\rho_{\rm NSC}(r)$ $=\frac{\bar{\rho}_{\rm NSC}}{1+\left(\frac{r}{0.22{\rm pc}}\right)^{n_{\rm NSC}}},~~r\leq 200{\rm pc},$ (2)
where $r=\sqrt{R^{2}+z^{2}}$, $\bar{\rho}_{\rm NSD}=330M_{\odot}/{\rm pc}^{3}$, $\bar{\rho}_{\rm NSC}=3.3\times 10^{6}M_{\odot}/{\rm pc}^{3}$ and $n_{\rm NSC}=2$ for $r<6{\rm pc}$ and $\bar{\rho}_{\rm NSC}=9.0\times 10^{7}M_{\odot}/{\rm pc}^{3}$ and $n_{\rm NSC}=3$ for $r\geq 6{\rm pc}$.
For this cylindrically symmetric baryon density profile, the waveform features an ansatz of $\psi(t,R,z)=e^{-i\gamma t/\hbar}\phi(R,z)$, where $\gamma$ is the ansatz energy eigenvalue. Then the FDM soliton density $\rho(R,z)=|\psi|^{2}=\phi^{2}(R,z)$ is simply related to the FDM soliton mass $M=\int_{R=0}^{R=\infty}\int_{z=-\infty}^{z=+\infty}2\pi R\rho(R,z)dRdz$, hence the total baryon mass inside the FDM soliton $M_{\rm b}=\int_{R=0}^{R=\infty}\int_{z=-\infty}^{z=+\infty}2\pi R\rho_{\rm b}(R,z)dRdz$. Also with this cylindrical ansatz, Eq. (1) can be simplified to
$\begin{cases}\begin{aligned} &\tilde{\nabla}^{2}\tilde{\phi}=2(\tilde{\Phi}+\tilde{\gamma})\tilde{\phi},\\ &\tilde{\nabla}^{2}\tilde{\Phi}=\tilde{\phi}^{2}+\tilde{\rho}_{\rm b},\end{aligned}\end{cases}$ (3)
where a set of dimensionless variables is defined as
$\tilde{\phi}\equiv\frac{\hbar\sqrt{4\pi G}}{mc^{2}}\phi,~~\tilde{\rho}_{\rm b}\equiv\left(\frac{\hbar\sqrt{4\pi G}}{mc^{2}}\right)^{2}\rho_{\rm b},$
$\tilde{M}\equiv\frac{GMm}{\hbar c},~~\tilde{M}_{\rm b}\equiv\frac{GM_{\rm b}m}{\hbar c},$
$\tilde{r}\equiv\frac{mc}{\hbar}r,~~\tilde{\Phi}\equiv\frac{1}{c^{2}}\Phi,~~\tilde{\gamma}\equiv\frac{1}{mc^{2}}\gamma.$ (4)
This dimensionless SP system obeys the scaling symmetry when
$\tilde{\phi}\longrightarrow\lambda\tilde{\phi},~~\tilde{\rho}_{\rm b}\longrightarrow\lambda^{2}\tilde{\rho}_{\rm b},$
$\tilde{M}\longrightarrow\lambda^{1/2}\tilde{M},~~\tilde{M}_{\rm b}\longrightarrow\lambda^{1/2}\tilde{M}_{\rm b}$
$\tilde{r}\longrightarrow\lambda^{-1/2}\tilde{r},~~\tilde{\Phi}\longrightarrow\lambda\tilde{\Phi},~~\tilde{\gamma}\longrightarrow\lambda\tilde{\gamma}$. (5)
[…]
$M\approx 1.35\times 10^{9}\left(\frac{M_{\rm{halo}}}{10^{12}M_{\odot}}\right)^{1/3}\left(\frac{m}{10^{-22}~{\rm{eV}}/c^{2}}\right)^{-1}M_{\odot}.$ (6)
When $\lambda_{0}=\lambda_{1}$, it means that $\rho_{\rm b}$ and $M$ are equal to the baryon density and the FDM soliton mass of the Milky Way at the same time. Therefore, if $\lambda_{0}\neq\lambda_{1}$, we replace $\lambda_{0}$ with $\lambda_{1}$ and repeat the above three steps Tan:2024dne . We stop the iteration until $\lambda_{1}/\lambda_{0}-1>0.0001$ or $\lambda_{0}\approx\lambda_{1}\approx 6\times 10^{-7}$ for [^1] Here $m=10^{-22}~\rm{eV}/c^{2}$ just serve as an example for didactic purpose. Although different $m$ leads to different results in the following discussion, the method proposed in this paper could still hold. $m=10^{-22}~\rm{eV}/c^{2}$. In Fig. 1 and Fig. 2, we show the final solutions of $\rho_{\rm b}$ and $\Phi$, which are consistent with the results of  Bar:2019bqz . These solutions are dependent on $\tan\theta=\frac{z}{R}$, which will serve as mock data to construct the baryon EoM inside the FDM soliton.
Figure 1: FDM soliton density contours in the inner Milky Way, which is consistent with Fig. 2 of Bar:2019bqz . The corresponding dimensionless total baryon mass is $\tilde{M}_{\rm b}=1.03$.
Figure 2: Total potential contours inside the FDM soliton, where $\tilde{\Phi}(\tilde{R}=0,\tilde{z}=0)=-2.75395$ which is accompanied by $\tilde{\rho}_{\rm b}(\tilde{R}=0,\tilde{z}=0)=21551.2$ and $\tilde{\phi}^{2}(\tilde{R}=0,\tilde{z}=0)=1$.
III Construction of baryon EoM by machine learning
GA as a machine learning technique is inspired by the concepts in evolutionary biology: a population evolves generation by generation to adapt to the environment; a population is composed of many individuals; each individual is characterized by its own chromosome; several genes form the chromosome; every gene includes some nucleotides. Therefore, making an analogy with the individual taking the form of
${\rm individual}=\sum^{\rm chromosome}{\rm gene}({\rm nucleotides}),$ (7)
in this section, we look for a mathematical expression optimizing the goodness of fit to the given data of $\tilde{\rho}_{\rm b}$. First, we describe how the mock data of $\tilde{\rho}_{\rm b}$ for the fitting process was gathered. Then we construct the fitting function or analytical emulator for baryon EoM by GA.
III.1 Mock data
Since Fig. 1 shows the distribution of $\rho(R,z)$, from it and by setting $\lambda\approx 6\times 10^{-7}$, we can extract $\tilde{\phi}^{2}(\tilde{R},\theta)$ for a given pair of $\{\tilde{R},\theta\leq 45^{\circ}\}$ or $\tilde{\phi}^{2}(\tilde{z},\theta)$ for a given pair of $\{\tilde{z},\theta>45^{\circ}\}$. Here, we consider that $\theta=0^{\circ},1^{\circ},...,90^{\circ}$ and $\Delta\tilde{R}=\Delta\tilde{z}=0.01$. However, we only sample $91\times 91$ $\tilde{\phi}^{2}$ points by $91\times 46$ pairs of $\{\tilde{R},\theta\leq 45^{\circ}\}$ and $91\times 45$ pairs of $\{\tilde{z},\theta>45^{\circ}\}$. Similarly, we also sample $91\times 91$ $\tilde{\Phi}$ points by the same pairs of $\{\tilde{R},\theta\leq 45^{\circ}\}$ and $\{\tilde{z},\theta>45^{\circ}\}$ from Fig. 2. Combing the corresponding $91\times 91$ $\tilde{\rho}_{\rm b}$ points retrieved from Eq. (II), our final dataset is a table of dimensions $8281\times 4$ with data points given as $\{\theta,\tilde{\phi}^{2},\tilde{\Phi},\tilde{\rho}_{\rm b}\}$. We quantify the goodness of fit of a given analytical expression for baryon EoM by the following function
$\%{\rm Acc}=\frac{100}{N}\sum_{i}^{N}\left|\frac{\lg[\tilde{\rho}_{{\rm b},i}]-\lg[\tilde{\rho}_{{\rm b},i,{\rm EoM}}]}{\lg[\tilde{\rho}_{{\rm b},i}]}\right|$ (8)
where $N=91\times 91=8281$ is the number of data points in our dataset and $\tilde{\rho}_{{\rm b},i}$ is a simplified notation for $\tilde{\rho}_{\rm b}$ evaluated at $\{\theta_{i},\tilde{\phi}^{2}_{i},\tilde{\Phi}_{i}\}$.
III.2 Analytical emulator for baryon EoM
GA looks for a fitting formula to a dataset by evolving the population, each individual of which takes a form similar to Eq. (7). More precisely, in this paper, we set the analytical emulator for baryon EoM as
$\lg[\tilde{\rho}_{\rm b,EoM}(x,\theta)]$ $\equiv\lg 21551.2+f_{\rm EoM}(3.75395,\theta)-f_{\rm EoM}(x,\theta),$
$f_{\rm EoM}(x,\theta)$ $\equiv\sum_{i=1}^{7}a_{i}(\theta)g_{i}\{b_{i}(\theta)[x-c_{i}(\theta)]\}d_{i}(x),$
$x$ $\equiv\tilde{\phi}^{2}-\tilde{\Phi}$, (9)
where $21551.2$ comes from the maximum value of the mock data of $\tilde{\rho}_{\rm b}$. The sum goes from 1 to 7 because the chromosome of each individual contains 7 genes and each gene is composed of several nucleotides $\{a_{i},b_{i},c_{i},d_{i},g_{i}\}$. $g_{i}$ is an operation in the grammar set $\{x,e^{x},-e^{-x}\}$ and, for simplicity, we fix them as
$g_{1}(x)$ $=g_{2}(x)=e^{x},$
$g_{3}(x)$ $=g_{4}(x)=g_{5}(x)=g_{6}(x)=-e^{-x},$
$g_{7}(x)$ $=-x.$ (10)
$\{a_{i},b_{i},c_{i}\}$ are linear functions of $\sin\theta$ to characterize the cylindrical symmetry of $f_{\rm EoM}(x,\theta)$
$a_{i}(\theta)$ $=a^{\prime}_{i}(1+a^{\prime\prime}_{i}\sin\theta),$
$b_{i}(\theta)$ $=b^{\prime}_{i}(1+b^{\prime\prime}_{i}\sin\theta),$
$c_{i}(\theta)$ $=c^{\prime}_{i}(1+c^{\prime\prime}_{i}\sin\theta),$ (11)
where $\{a^{\prime}_{i},b^{\prime}_{i},c^{\prime}_{i}\}\in[0,5]$ and $\{a^{\prime\prime}_{i},b^{\prime\prime}_{i},c^{\prime\prime}_{i}\}\in[-1,1]$ are random numbers. $d_{i}$ is a correction that can be fine-tuned to improve the goodness of fit
$d_{i}(x)=\{1+0.18\left|j_{1}[1.5\pi(x-3.75395)]\right|\}(\delta_{1i}+\delta_{2i}),\\$ (12)
where $\delta_{1i}$ and $\delta_{2i}$ mean that this correction only works for $i=1$ and $i=2$ cases and $j_{1}(x)=(\sin x-x\cos x)/x^{2}$ is the spherical Bessel function of order 1.
We perform a stochastic search with GA code Arjona:2019fwb ; Orjuela-Quintana:2022nnq ; Orjuela-Quintana:2023uqb ; Orjuela-Quintana:2024hha and find a function with $\%{\rm Acc}=2.54844$, whose parameters are listed in Tab. 1. We also plot the accuracy of this analytical emulator for baryon EoM at different $\theta$ in Fig. 3. The accuracy is $\%4$, which means that our emulator can be used for some simple baryon evolutions inside the FDM soliton. In fact, the accuracy at $\theta\gtrsim 60^{\circ}$ can be further improved if the unphysical jump in mock data $\tilde{\rho}_{\rm b}$ due to the cutoff of Eq. (II) disappears. In Fig. 4, the comparison between $\tilde{\rho}_{\rm b,EoM}(x,\theta)$ (colored points) and $\tilde{\rho}_{\rm b}(x,\theta)$ (black points) at four angles $\theta=\{0^{\circ},30^{\circ},60^{\circ},90^{\circ}\}$ is shown respectively.
Table 1: Best-fit parameters for the baryon EoM emulator given in Eq. (III.2).
$a^{\prime}_{i}$ $a^{\prime\prime}_{i}$ $b^{\prime}_{i}$ $b^{\prime\prime}_{i}$ $c^{\prime}_{i}$ $c^{\prime\prime}_{i}$
$i=1$ 1.35941 -0.31828 0.75068 -0.02252 0.45798 0.04361
$i=2$ 1.11927 0.39288 0.19257 -0.90871 2.99997 -0.90871
$i=3$ 1.11756 0.22907 0.56427 -0.56092 2.92398 -0.48616
$i=4$ 3.83729 0.54773 0.89039 -0.15683 0.73105 -0.72978
$i=5$ 0.05903 0.88515 3.49499 0.99812 1.37196 -0.41290
$i=6$ 0.31317 0.99983 3.88515 -0.07250 1.01848 0.02457
$i=7$ 2.00000 -0.20860 2.88828 -0.61984 1.52037 -0.72819
Figure 3: Accuracy of the analytical emulator for baryon EoM as a function of $\theta$.
Figure 4: Comparison between the analytical emulator for baryon EoM $\tilde{\rho}_{\rm b,EoM}(x,\theta)$ (colored points) and the mock data $\tilde{\rho}_{\rm b}(x,\theta)$ (black points) at four angles $\theta=\{0^{\circ},30^{\circ},60^{\circ},90^{\circ}\}$ respectively. The unphysical jump in $\tilde{\rho}_{\rm b}(x,\theta)$ for larger $\theta$ comes from the cutoff of Eq. (II), which decreases the accuracy to some extent.
IV FDM soliton formation with an analytical baryon EoM
Note that the above analytical emulator for baryon EoM is constructed from dimensionless mock data $\{\theta,\tilde{\phi}^{2},\tilde{\Phi},\tilde{\rho}_{\rm b}\}$. Therefore, replacing the fixed baryon background $\tilde{\rho}_{\rm b}$ of Eq. (3) with an analytical baryon EoM $\tilde{\rho}_{\rm b,EoM}$ (Eq. (III.2)) leads to a new coupled system
$\begin{cases}\begin{aligned} &\tilde{\nabla}^{2}\tilde{\phi}=2(\tilde{\Phi}+\tilde{\gamma})\tilde{\phi},\\ &\tilde{\nabla}^{2}\tilde{\Phi}=\tilde{\phi}^{2}+\tilde{\rho}_{\rm b,EoM},\\ &\lg\tilde{\rho}_{\rm b,EoM}=4.3+f_{\rm EoM}(3.75,\theta)-f_{\rm EoM}(x,\theta),\end{aligned}\end{cases}$ (13)
[…]
$\delta\rho(R,z)$ $=\frac{\phi^{2}_{\rm Eq13}(R,z)-\phi^{2}_{\rm Eq3}(R,z)}{\phi^{2}_{\rm Eq3}(R,z)},$ (14)
$\delta\Phi(R,z)$ $=\frac{\Phi_{\rm Eq13}(R,z)-\Phi_{\rm Eq3}(R,z)}{\Phi_{\rm Eq3}(R,z)}.$ (15)
We find both of $|\delta\rho(R,z)|$ and $|\delta\Phi(R,z)|$ are $\lesssim 0.04$, which is consistent with the accuracy of our analytical emulator for baryon EoM $\%4$, as shown in Fig. 3.
Figure 5: Distribution of fractional error of $\delta\rho(R,z)$.
Figure 6: Distribution of fractional error of $\delta\Phi(R,z)$.
V Summary and discussion
In this paper, we first solve the SP system with a fixed empirical baryon density profile background in cylindrical coordinates by an iteration of $\lambda$. For $m=10^{-22}\rm{eV}/c^{2}$, $\lambda\approx 6\times 10^{-7}$ rescales the dimensionless $\tilde{\rho}_{\rm b}$ and $\tilde{M}$ to the baryon density profile and the FDM soliton mass of the Milky Way simultaneously. Therefore, $\lambda\approx 6\times 10^{-7}$ was used to map the scaled dimensionless variables to the corresponding physical dimensional ones and vice versa. Then we make the obtained stationary solutions mock data and construct an analytical emulator (with $\%{\rm Acc}<\%4$) for the baryon EoM inside the FDM soliton by GA, using a machine learning technique. Finally, we complement a model validation by comparison between the stationary solutions of an enlarged SP system with the emulator $\tilde{\rho}_{\rm b,EoM}$ as the third equation under certain conditions and ones of the original SP system. We find the fractional errors of two sets of solutions are $\lesssim 0.04$ and are consistent with the accuracy of emulator for baryon EoM $\%4$.
What about its validation for the nonstationary evolutions of FDM solitons, especially those evolutions totally departing from the stationary solutions, such as FDM soliton collisions. Since the mock data used in this paper is extracted from the stationary solutions of the SP system with a fixed empirical baryon density profile background, our analytical emulator for the baryon EoM inside the FDM soliton should also be valid for the perturbation, the tidal disruption or deformation of FDM solitons conservatively. However, for some dramatic evolutions, our emulator probably does not perform correctly.
Acknowledgments We acknowledge the use of HPC Cluster of Tianhe II in National Supercomputing Center in Guangzhou. Ke Wang is supported by grant from the National Key Research and Development Program of China (grant No. 2021YFC2203003). Jianbo Lu is supported by the National Natural Science Foundation of China (12175095), and supported by LiaoNing Revitalization Talents Program (XLYC2007047). Man Ho Chan is supported by the grant from the Research Grants Council of the Hong Kong Special Administrative Region, China (Project No. EdUHK 18300324).